# IPS function for the Schrödinger equation

### v1

@@ 1. Introduction
in $H^{1}_{\text{loc}}(\Omega\setminus\sigma)$ (see also Proposition 1 in [4] for its origin). Then, a similar argument for the derivation of (1.3) one gets also
@@ 1.1. IPS and Side A of Singular Sources Method
modulo $O(1)$.
@@ 1.3. Comparison to Singular Sources Method of Potthast
$\bullet$ The choice (1.27) avoids a technical difficulty caused by the presence of polarization. One obtains the blowing-up property of $\nabla\cdot\mbox{$w$}_{x}(x)$ as $x$ simply approaches point $a\in\partial D$ without considering polarization. Besides, by using (1.16) and Remark 1.1, one obtain the lower estimate of the bowing-up rate as $x\rightarrow a$
@@
This paper is organized as follows. Theorem 1.2 is proved in Subsection 2.1 as a special case of Proposition 2.1. It is derived by using Green’s representation formula of solutions of (1.5) and (1.7), that is, $\mbox{$w$}_{x}$ and $\mbox{$w$}_{x}^{1}$. This proposition introduces a decomposition formula of $\nabla\cdot\mbox{$W$}_{x}(y)$ for all $(x,y)\in(\Omega\setminus\overline{D})^{2}$ in terms of the functions ${\mathcal{I}}(x,y)$ and ${\mathcal{I}}^{1}(x,y)$ defined by (2.2) and (2.3), which we call the liftings of ${\mathcal{I}}(x)$ and ${\mathcal{I}}^{1}(x)$, respectively. Subsection 2.2 lists and proves their properties, especially the energy integral representation, symmetry, and harmonicity when one of the two variables $x$ and $y$ is fixed. Theorems 1.3 and 1.4 are proved in Subsections 2.3 and 2.4, respectively. In Subsection 2.5, as an advantage of introducing the lifting of ${\mathcal{I}}(x)$, we describe Theorem 2.1 which states that the local proving in an arbitrary (small) neighbourhood of a given point on $\partial\Omega$ relative to $\overline{\Omega}$ uniquely determines the indicator function ${\mathcal{I}}(x)$ in the domain $\Omega\setminus\overline{D}$ globally. Section 3 is devoted to Side B of IPS, which states the blowing up property of three types of sequences calculated from Dirichlet-to-Neumann map. They are called the indicator sequences and in Theorem 3.1, under stronger assumption on the jump of $V$, their blowing up are proved. The key for the proof is Proposition 3.2, which states, roughly speaking, on an open set having intersection with a needle, the corresponding needle sequence blows up and its $L^{1}$-norm is weaker than $L^{2}$-norm. This last property is a new fact. As a consequence of Theorem 3.1 a characterization of obstacle is derived as Corollaries 3.1 to 3.3. The final section describes some remarks on problems for future consideration. In the Appendix, we provide the proof of Lemma A for the reader’s convenience. This lemma is needed for Proposition 3.1 of Section 3, which concernes the blowing up property of needle sequences.
@@ 2.5. Uniqueness from local probing
Let $U$ and $V$ be nonempty open subsets of $\mathbb{R}^{3}$ such that $U\subset\Omega$ and $\overline{U}\cap\overline{D}=\emptyset$. Then the values of ${\mathcal{I}}(x,y)$ for all $(x,y)\in(\Omega\setminus\overline{D})^{2}$ and thus ${\mathcal{I}}(x)={\mathcal{I}}(x,x)$ for all $x\in\Omega\setminus\overline{D}$ are uniquely determined by those of ${\mathcal{I}}(x,y)$ for all $(x,y)\in U\times V$.

### v2

@@ 1. Introduction
in $H^{1}_{\text{loc}}(\Omega\setminus\sigma)$ (see also Proposition 1 in [4] for its origin). Then, a similar argument for the derivation of (1.3) yields also
@@ 1.1. IPS and Side A of Singular Sources Method
modulo $O(1)$. Thus it follows from (b) of Theorem 1.1 that, for each $a\in\partial D$ we have
[…]
Furthermore it holds that, for each point $b\in\partial\Omega$
$\lim_{x\rightarrow b}\nabla\cdot\mbox{$W$}_{x}(x)=-\infty.$
This needs some technical argument and for the proof see Appendix.
@@ 1.3. Comparison to Singular Sources Method of Potthast
$\bullet$ The choice (1.27) avoids a technical difficulty caused by the presence of polarization. One obtains the blowing-up property of $\nabla\cdot\mbox{$w$}_{x}(x)$ as $x$ simply approaches point $a\in\partial D$ without considering polarization. Besides, by using (1.16) and Remark 1.1, one obtains the lower estimate of the bowing-up rate as $x\rightarrow a$
@@
This paper is organized as follows. Theorem 1.2 is proved in Subsection 2.1 as a special case of Proposition 2.1. It is derived by using Green’s representation formula of solutions of (1.5) and (1.7), that is, $\mbox{$w$}_{x}$ and $\mbox{$w$}_{x}^{1}$. This proposition introduces a decomposition formula of $\nabla\cdot\mbox{$W$}_{x}(y)$ for all $(x,y)\in(\Omega\setminus\overline{D})^{2}$ in terms of the functions ${\mathcal{I}}(x,y)$ and ${\mathcal{I}}^{1}(x,y)$ defined by (2.2) and (2.3), which we call the liftings of ${\mathcal{I}}(x)$ and ${\mathcal{I}}^{1}(x)$, respectively. Subsection 2.2 lists and proves their properties, especially the energy integral representation, symmetry, and harmonicity when one of the two variables $x$ and $y$ is fixed. Theorems 1.3 and 1.4 are proved in Subsections 2.3 and 2.4, respectively. In Subsection 2.5, as an advantage of introducing the lifting of ${\mathcal{I}}(x)$, we describe Theorem 2.1 which states that the local proving in an arbitrary (small) neighbourhood of a given point on $\partial\Omega$ relative to $\overline{\Omega}$ uniquely determines the indicator function ${\mathcal{I}}(x)$ in the domain $\Omega\setminus\overline{D}$ globally. Section 3 is devoted to Side B of IPS, which states the blowing up property of three types of sequences calculated from Dirichlet-to-Neumann map. They are called the indicator sequences and in Theorem 3.1, under stronger assumption on the jump of $V$, their blowing up are proved. The key for the proof is Proposition 3.2, which states, roughly speaking, on an open set having intersection with a needle, the corresponding needle sequence blows up and its $L^{1}$-norm is weaker than $L^{2}$-norm. This last property is a new fact. As a consequence of Theorem 3.1 a characterization of obstacle is derived as Corollaries 3.1 to 3.3. The final section describes some remarks on problems for future consideration. In the Appendix, first we provide the proof of Lemma A for the reader’s convenience. This lemma is needed for Proposition 3.1 of Section 3, which concernes the blowing up property of needle sequences. The second is devoted to a proof of the blowing-up of the IPS function on $\partial\Omega$.
@@ 2.5. Uniqueness from local probing
Let $U$ and $V$ be nonempty open subsets of $\mathbb{R}^{3}$ such that $U\cup V\subset\Omega$ and $\overline{U}\cap\overline{D}=\overline{V}\cap\overline{D}=\emptyset$. Then the values of ${\mathcal{I}}(x,y)$ for all $(x,y)\in(\Omega\setminus\overline{D})^{2}$ and thus ${\mathcal{I}}(x)={\mathcal{I}}(x,x)$ for all $x\in\Omega\setminus\overline{D}$ are uniquely determined by those of ${\mathcal{I}}(x,y)$ for all $(x,y)\in U\times V$.
@@ 5.1. Divergence of an integral
5.2. Blowing-up of the IPS function on $\partial\Omega$
In this subsection we give a proof of
Proposition B. Given $b\in\partial\Omega$, we have
$\lim_{x\rightarrow b}\,{\mathcal{I}}^{1}(x)=-\infty.$
Once we have this, then the following statement is deduced from (c) of Theorems 1.1 and 1.2: it holds that, for each $b\in\partial\Omega$
$\lim_{x\rightarrow b}\nabla\cdot\mbox{$W$}_{x}(x)=-\infty.$
Proof of Proposition B is based on a crucial estimate.
Lemma B.1. Let $W$ be a bounded domain with $C^{1,1}$-boundary of $\mathbb{R}^{3}$. Let $V\in L^{\infty}(W)$. Let $v\in H^{1}(W)$ satisfy $\Delta v+V(z)v=0$ in $W$. Assume that, 0 is not a Dirichlet eigenvalue of $\Delta+V(z)$ in $W$. Then, there exists a positive constant $C$ independent of $v$ such that
$\|v\|_{L^{2}(W)}^{2}\leq C\|v|_{\partial W}\|_{L^{2}(\partial W)}^{2}.$ (B.1)
For a proof, apply the argument for the proof of (A.2) of Proposition A.1 in [6].
Proof of Proposition B. Applying (B.1) to the solution of (1.7), one gets
$\|\mbox{$w$}_{x}^{1}\|_{L^{2}(\Omega)}^{2}\leq C\|\nabla G(\,\cdot\,-x)|_{\partial\Omega}\|_{L^{2}(\partial\Omega)}^{2}.$
Then, by (1.10) we obtain
$-{\mathcal{I}}^{1}(x)\geq\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}-C^{\prime}\|\nabla G(\,\cdot\,-x)|_{\partial\Omega}\|_{L^{2}(\partial\Omega)}^{2},$
where $C^{\prime}$ is a positive constant. Thus it suffices to prove that
$\lim_{x\rightarrow b}\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}=\infty$ (B.2)
and
$\lim_{x\rightarrow b}\frac{\|\nabla G(\,\cdot\,-x)|_{\partial\Omega}\|_{L^{2}(\partial\Omega)}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}=0.$ (B.3)
Ignoring a constant multiplication, we have
$\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}\sim\int_{\mathbb{R}^{3}\setminus\overline{\Omega}}\frac{dz}{|z-x|^{6}}.$
Then from Lemma B.2 mentioned below we obtain (B.2) and (B.3).
$\Box$
Lemma B.2. Assume that $\partial\Omega$ be Lipschitz. We have, as $x\rightarrow b$ in $\Omega$
$\int_{\mathbb{R}^{3}\setminus\overline{\Omega}}\frac{dz}{|z-x|^{6}}\geq\frac{C_{1}}{|x-b|^{3}}$ (B.4)
and
$\frac{\|\nabla G(\,\cdot\,-x)|_{\partial\Omega}\|_{L^{2}(\partial\Omega)}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}\leq C_{2}\sqrt{|x-b|},$ (B.5)
where $C_{1}$ and $C_{2}$ are positive constants.
Proof. We follow the proof of Theorem C in [4]. Choose the local coordinates at the point $b$ in such a way that
$(\mathbb{R}^{3}\setminus\overline{\Omega})\cap B(b,\delta)=\left\{z=b+A\left(\begin{array}[]{l}\displaystyle s\\ \\ \displaystyle t\end{array}\right)\,\left|\right.\,\varphi(s)<t\,,|s|^{2}+t^{2}<\delta^{2}\right\},$
where $\delta=\delta(b)$ is a small positive constant, $(s,t)\in\mathbb{R}^{2}\times\mathbb{R}$, $\varphi\in C^{0,1}(\mathbb{R}^{2})$ with compact support and satisfies $\varphi(0)=0$, and $A$ is a $3\times 3$-orthogonal matrix.
First of all we have
$\begin{array}[]{ll}\displaystyle\int_{\mathbb{R}^{3}\setminus\overline{\Omega}}\frac{dz}{|z-x|^{6}}&\displaystyle\geq\int_{(\mathbb{R}^{3}\setminus\overline{\Omega})\cap B(b,\delta_{b})}\frac{dz}{|z-x|^{6}}.\end{array}$ (B.6)
Since
$|z-x|\leq(|s|^{2}+t^{2})^{\frac{1}{2}}+|x-b|,$
one gets
$\begin{array}[]{ll}\displaystyle\int_{(\mathbb{R}^{3}\setminus\overline{\Omega})\cap B(b,\delta_{b})}\frac{dz}{|z-x|^{6}}&\displaystyle\geq C\int_{\varphi(s)<t,\,|s|^{2}+t^{2}<\delta^{2}}\frac{dsdt}{\{(|s|^{2}+t^{2})^{\frac{1}{2}}+|x-b|\}^{6}}\\ \\ \displaystyle&\displaystyle=\frac{C}{|x-b|^{3}}\int_{\varphi(|x-b|s)<|x-b|t,\,|s|^{2}+t^{2}<\delta^{2}|x-b|^{-2}}\frac{dsdt}{\{(|s|^{2}+t^{2})^{\frac{1}{2}}+1\}^{6}}.\end{array}$ (B.7)
Here we use the change of variables
$\begin{array}[]{ll}\displaystyle\left(\begin{array}[]{l}\displaystyle s\\ \\ \displaystyle t\end{array}\right)=\rho\,\mbox{$\omega$},&\rho>0,\,\mbox{$\omega$}=(\omega^{\prime},\omega_{3})\in S^{2},\omega^{\prime}=(\omega_{1},\omega_{2}).\end{array}$
One gets
$\begin{array}[]{l}\,\,\,\,\,\,\displaystyle\int_{\varphi(|x-b|s)<|x-b|t,\,|s|^{2}+t^{2}<\delta^{2}|x-b|^{-2}}\frac{dsdt}{\{(|s|^{2}+t^{2})^{\frac{1}{2}}+1\}^{6}}\\ \\ \displaystyle=\int_{S^{2}}\int_{\varphi(|x-b|\rho\omega^{\prime})<|x-b|\rho\omega_{3},\,\rho^{2}<\delta^{2}|x-b|^{-2}}\frac{\rho^{2}d\rho d\mbox{$\omega$}}{(\rho+1)^{6}}.\end{array}$ (B.8)
Choose a positive constant $K$ such that $|\varphi(s)|\leq K|s|$. Let $\omega$ satisfies $\omega_{3}>K|\omega^{\prime}|$. Then we have
$\begin{array}[]{ll}\displaystyle\varphi(|x-b|\rho\omega^{\prime})&\displaystyle\leq K|x-b|\rho|\omega^{\prime}|\\ \\ \displaystyle&\displaystyle<|x-b|\rho\omega_{3}.\end{array}$
Thus one gets
$\begin{array}[]{l}\displaystyle\int_{S^{2}}\int_{\varphi(|x-b|\rho\omega^{\prime})<|x-b|\rho\omega_{3},\,\rho^{2}<\delta^{2}|x-b|^{-2}}\frac{\rho^{2}d\rho\,d\mbox{$\omega$}}{(\rho+1)^{6}}\\ \\ \displaystyle\geq\int_{\omega\in S^{2},\,\omega_{3}>K|\omega^{\prime}|}d\mbox{$\omega$}\int_{0}^{\frac{\delta}{|x-b|}}\frac{\rho^{2}}{(\rho+1)^{6}}d\rho.\end{array}$ (B.9)
Since
$\int_{\omega\in S^{2},\,\omega_{3}>K|\omega^{\prime}|}d\mbox{$\omega$}>0$
and
$0<\int_{0}^{\infty}\frac{\rho^{2}}{(\rho+1)^{6}}d\rho<\infty,$
one obtains, as $x\rightarrow b$
$\int_{\omega\in S^{2},\,\omega_{3}>K|\omega^{\prime}|}d\mbox{$\omega$}\int_{0}^{\frac{\delta}{|x-b|}}\frac{\rho^{2}}{(\rho+1)^{6}}d\rho\geq C^{\prime},$
where $C^{\prime}$ is a positive constant. Therefore, from this together with (B.6) to (B.9), we obtain (B.4).
The proof of (B.5) is as follows. By Theorem 1.5.1.10 in [3], taking a large open ball $B$, we have, for all $\epsilon\in\,]0,\,1[$
$\|\nabla G(\,\cdot\,-x)\|_{L^{2}(\partial\Omega)}^{2}\leq K(\epsilon\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega})}^{2}+\epsilon^{-1}\|\nabla G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega})}^{2}),$
where $K$ is a positive constant independent of $\epsilon$ and $x\in\Omega$. Thus one gets
$\frac{\|\nabla G(\,\cdot\,-x)|_{\partial\Omega}\|_{L^{2}(\partial\Omega)}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}\leq K\left(\epsilon+\epsilon^{-1}\frac{\|\nabla G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega)}}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}\right).$ (B.10)
Here we have, with the help of Hölder’s inequality
$\begin{array}[]{ll}\displaystyle\|\nabla G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega)}}^{2}&\displaystyle=C_{1}\int_{B\setminus\overline{\Omega}}\frac{dz}{|z-x|^{4}}\\ \\ \displaystyle&\displaystyle\leq C_{2}\left(\int_{B\setminus\overline{\Omega}}\frac{dz}{|z-x|^{6}}\right)^{\frac{2}{3}}\\ \\ \displaystyle&\displaystyle=C_{3}\left(\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega})}^{2}\right)^{\frac{2}{3}}\\ \\ \displaystyle&\displaystyle=C_{3}\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega})}^{\frac{4}{3}},\end{array}$
where $C_{i}$, $i=1,2,3$ are positive constants. This yields
$\frac{\|\nabla G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega)}}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}\leq\frac{C}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{\frac{2}{3}}}.$
From this together with (B.4) one has, as $x\rightarrow b$
$\frac{\|\nabla G(\,\cdot\,-x)\|_{L^{2}(B\setminus\overline{\Omega)}}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}\leq C|x-b|.$
Thus from (B.10) one gets
$\frac{\|\nabla G(\,\cdot\,-x)|_{\partial\Omega}\|_{L^{2}(\partial\Omega)}^{2}}{\|\nabla^{2}G(\,\cdot\,-x)\|_{L^{2}(\mathbb{R}^{3}\setminus\overline{\Omega})}^{2}}\leq K(\epsilon+\epsilon^{-1}C|x-b|).$
Thus, as $x\rightarrow b$, choosing a small $\epsilon$ in such a way that
$\epsilon=\epsilon^{-1}C|x-b|,$
we obtain (B.5).
$\Box$
Note that the proof of (B.5) presented above never makes use of any upper bound of $L^{2}$-norm of $\nabla G(\,\cdot\,-x)$ over $\partial\Omega$.